OpenStudy (samsan9):

Solve: x^2 -6x + 58 = 0 A.) {3+7i, 3-7i} B.) {2+9i, 2-9i} C.) {7+3i, 7-3i} D.) {2+19i, 2-19i}

OpenStudy (zzr0ck3r):

\(x^2-6x+58=0\iff(x-3)^2=-58+(-3)^2=-58+9=-49\) so \((x-3)^2=-49\) can you complete it>

OpenStudy (zzr0ck3r):

?

OpenStudy (zzr0ck3r):

you can see the answer from there if you know what \(\sqrt{-49}=\)

OpenStudy (samsan9):

7i?

OpenStudy (zzr0ck3r):

right, so which one do you think it is?

OpenStudy (samsan9):

A

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

I completed the square, google it if you dont know it. Its not that hard and helps when there is no coefficient on the 1st term.